OpenStudy (anonymous):

The supplement of an angle is 4 more than twice its complement. Find the measure of the angle. is it 184?

OpenStudy (anonymous):

let angle be x 180 - x = 2(90-x) + 4 180 - x = 180 - 2x + 4 x = 4

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

ok - plz click on good answer to close the question